Prove(trigonometry):
tan(15)tan(25)tan(35)=tan(5)
Thanks

Grade:9

1 Answers

Aditya Gupta
5 years ago
so you have to use the formula: tanx*tan(60 – x)*tan(60+x)= tan3x
put x = 25 we get
tan25tan35tan85=tan75
using tan85=cot5=1/tan5 and tan75=1/tan15 we have
tan15tan25tan35= tan5